# Automatic Depression Assessment using Machine Learning: A Comprehensive Survey

### v1

@@ 4.7.3. Audio-visual-language approaches
Hand-crafted approaches: Based on audio, visual and texts, Williamson et al. (Williamson et al., 2016) analyse both semantic contents and context from dialogue transcripts, where contents are modeled at the sentence-level via the GloVe embedding (Pennington et al., 2014) followed by PCA/ZCA to reduce dimension and sparse coding to extract final semantic features. To capture the joint dynamical facial behaviour properties during speech, a rank-ordered eigenspectrum is computed from multiple AUs, describing within-and cross-channel distributional properties of the multi-channel AU time series. Finally, the extracted audio (discussed in Sec. 4.5), visual and language features are individually fed to a Gaussian Staircase Model to make unimodal depression prediction, followed by a decision-level fusion. Building on a similar pipeline as AVEC 2013/2014 baselines, Pampouchidou et al. (Pampouchidou et al., 2016) achieve the clip-level multi-modal depression assessment via two strategies: (i) decision-level fusion of depression predictions made by clip-level visual representation (i.e., statistical and spectral features of facial landmarks, head motions and blinking rates), clip-level audio representation computed from various audio LLDs, and the clip-level language representation; and (ii) fusing multi-modal clip-level feature and then feeding it to a SVR or RF for ADA. Gupta et al. (Gupta et al., 2014) extract frame-level LBP features and facial landmarks, alongside LBP-TOP and optical-flow-based motion vectors to capture short-term facial dynamics, all of which summarised as clip-level statistics. Meanwhile, multiple audio LLDs, spectral shape, spectro-temporal modulations, pitch periodicity, and long-term spectral variability, are jointly form a clip-level audio representation. Linguistic features are derived by mapping word sequences to affective ratings, processed via a Distributional Semantics Model and PCA as clip-level representation. The three modalities are then integrated and fed into an SVR for clip-level ADA. Nasir et al. (Nasir et al., 2016) compute arithmetic means of: (1) frame-level pitch, glottal and voice quality-related as well as vocal tract audio (e.g., F1, F2 and F3) features; and (2) several facial behaviour primitives expressed in each 10s window. Additionally facial feature dynamics (i.e., velocity and acceleration), facial geometric and facial topology changes defined by facial landmarks are also computed. Finally, the feature-level fusion is achieved on clip-level audio, visual and language (statistics of phoneme rate and duration) representations, followed by RF for ADA.
@@
Alternatively, some multi-modal approaches learn visual depression cues from pre-extracted facial/visual behaviour primitives rather than original videos. Based on virtual assistant-guided sessions, Wu et al. (Wu et al., 2024) separately conduct naive decision-level fusion and feature-level fusion for various mid-level and low-level hand-crafted features extracted from text, audio, facial attributes, heart rate, and eye movement, where MLP is employed as the classifier for four-level depression classification. Yang et al. (Yang et al., 2018, 2016) first select several equal-length video segment within every video to balance the depressed/non-depressed training examples. Based on each segment, it learns a Histogram of Displacement Range representing the range and speed of facial landmark displacements. Meanwhile, segment-level statistics summarising hand-crafted LLD audio features are computed. Finally, segment-level audio-visual features are processed by a two-stream CNN to individually make depression predictions, which are fused with textual features via a decision tree. Gong et al. (Gong and Poellabauer, 2017) specifically explore the relationship between human audio, facial and language behaviours triggered by $8383$ different interview topics defined by DAIC-WOZ database (Gratch et al., 2014) and the corresponding depression severity, where language information are partially used for topic modelling. For each topic, clip-level statistics of frame-level audio LLDs and facial descriptors as well as LIWC-based textual features, are extracted. Finally, a vector summarising all topic-level multi-modal features is constructed to make the clip-level depression prediction.
@@
Deep learning approaches: CNN-RNN framework have been widely extended for audio-visual-language ADA. Zhang et al. (Zhang et al., 2024a) deep learn frame-level/short-term audio, facial and language features via three CNNs, and then model their long-term dynamics via a BiLSTM. Then, the features learned from three modalities are fused via an attention, followed by a MLP to make the depression prediction. Wei et al. (Wei et al., 2022) first apply CNN-BiLSTM to independently learn audio, visual and textual features, and then introduce a novel sub-attention to fuse these multi-modal clip-level features. Lin et al. (Yin et al., 2019) also propose a Hierarchy BiLSTM to process audio, visual and text data for ADA. Rodrigues et al. (Rodrigues Makiuchi et al., 2019) apply a GNN-LSTM to process deep spectrum audio features and deep face features provided by ResNet-50. It additionally employs BERT to extract language features from texts, followed by a CNN-LSTM model to learn clip-level language depression features. Finally, multi-modal features are concatenated and fed to a dense layer for ADA. The similar strategy was also employed by (Ray et al., 2019), where three BiLSTM networks are developed to obtain clip-level audio, visual and language representations from frame/segment-level hand-crafted and deep-learned features, followed by a multi-layer attention network to fuse them for ADA. In (Zhang et al., 2024c), a ResNet and Swin Transformer-based dual visual encoder first learns global and local spatial facial features from each frame, and weights each facial local region’s importance. A BiLSTM and a multi-head attention operation then model spatio-temporal depression cues from all frame-level representations. Meanwhile, the Mel spectrogram is calculated from the audio signal, followed by GRU and multi-head attention for temporal modelling. The text feature is modelled via a BERT Chinese model and attention. Finally, another attention module selects and combines depression features of three modalities, and a MLP is employed for the final prediction. Alternatively, some approaches achieve deep audio-visual-language ADA without using RNN-style networks, Haque et al. (Haque et al., 2018) extend the Causal CNN to learn sentence-level embeddings from audio, text and 3D facial landmarks, which are summarised into a single vector via simple average at the sentence-level, and then fed to SVM for ADA. Sun et al. (Sun et al., 2022b) propose two audio, visual and language fusion strategies for ADA, where one forms an cubeMLP based on three independent MLP units to mix multi-modal features across three axes, while the other (Sun et al., 2022a) designing a tensor-based multi-modal Transformer conducting source-target attentions between features of every modality pair. Zhang et al. (Zhang, [n. d.]) introduces two attention-based aggregation strategies to: (1) dynamically evaluate the relevance among audio, visual and language features to adaptively integrate their temporal dynamics; and (2) determine the importance of frame-level amalgamated features to select the most depression-specific features. To investigate dynamics of multi-modal behaviours, Shen et al. (Shen et al., 2024) propose a multi-modal behaviour graph representation learning framework which learns a set of heterogeneous node features to represent short segment-level facial behaviours, long segment-level audio behaviours, and the global text cues. Then, edges are constructed to connect nodes of the same and different modalities, to model multi-scale intra and inter-modal dynamics. This clip-level graph representation is then fed to a GNN for ADA.
4.7.4. Other multi-modal approaches
@@
A few studies also utilise other modality combination for ADA. Liu et al (Liu et al., 2022) propose a two-stream CNN for facial and body movement-based ADA, where one stream is pre-trained to predict frame-level facial expression, based on which a video-level expression entropy is computed. The other stream locates the position of the individual’s 18 body joints and calculate their video-level motion amplitude. Finally, a LR fuses video-level facial expression entropy and body joint motion amplitude for ADA. Ning et al. (Ning et al., 2024) propose an auxiliary decision-making system to combine EEG and speech signals for ADA. It explores both linear and nonlinear features from each modality, and concatenates them for depression classification. Li et al. (Li et al., 2024d) fuse a sparse emotion dictionary describing textual data and spectral audio features for ADA, where a multi-head self-attention-based low and high frequency feature fusion strategy is proposed. Ye et al. (Ye et al., 2021) extract several hand-crafted audio LLDs and apply DeepSpectrum to deep learn depression-related audio features, which are subsequently integrated into a customised TCN, with the final layer employing relational attention classification. Meanwhile, they employ the Continuous Bag of Words (CBOW) method for text feature extraction. Both clip-level features are concatenated and fed to a MLP for ADA. Zheng et al. (Zheng et al., 2023) propose a temporal convolution transformer to jointly assess depression and emotion status from human audio, visual, text and additional physiological behaviours. Fan et al. (Fan et al., 2024) leverage CNN, transformer and GNN to extract depression-related frame/segment-level video, audio and remote photoplethysmographic (rPPG) features and treats each of them as a graph node. Then, a transformer is introduced to model inter-modal, intra-modal, and tri-modal relationships among nodes learned from three modalities. Finally, a GNN fuses all features for clip-level multi-modal ADA. Lin et al. (Lin et al., 2021) extract various hand-crafted frame-level features from human body movements, facial behaviours and audio signals, which are summarised as a multi-modal clip-level representation through a Multi-modal Deep Denoising Auto-Encoder and Fisher vector encoding, followed by a RF for feature selection. Finally, MLP and LR are separately applied for ADA. Shao et al. (Shao et al., 2024a) propose to learn complementary depression-related cues from brain behaviours detected by EEG and fNIRS, where spatio-temporal CNNs and transformers are stacked to jointly process both modalities.
5. Publicly available automatic depression assessment datasets
@@
This section reviews existing publicly available ADA datasets. The main challenge is their limited scale in terms of participant number and clip length, i.e., they generally contain (except the V-log dataset that lacks rigorous labeling) fewer than 200 participants while each clip usually lasts less than 30 mins (i.e., depression is a long-term condition reflected over extended periods). Another limitation is that depression labels are typically obtained from self-reported questionnaires rather than standardised clinical assessments, leading to some degree of subjectivity and inconsistency.
Lack of large-scale and high-quality ADA datasets: Another common challenge is the lack of large-scale, high-quality and well-annotated publicly available datasets for researchers to develop and and fairly compare their multi-modal expressive behaviour-based ADA models, as existing public datasets (except the V-log dataset that lacks rigorous labeling) contain fewer than 200 participants. Depression labels are typically obtained from self-reported questionnaires rather than standardised clinical assessments, leading to some degree of subjectivity and inconsistency. Additionally, these publicly available datasets rely on short-term (less than 1 hour) recordings, despite depression being a long-term condition reflected over extended periods. Such issues not only limit the advancement of DL-based ADA models but also undermine evaluation objectivity. Alternatively, collecting a high-quality and large-scale depression dataset with diverse participant demographics and reliable depression labels remains challenging due to ethical, logistical, and financial constraints.
@@
The AVEC 2013 depression corpus (Valstar et al., 2013) is a pioneering publicly available audio-visual ADA dataset containing 150 naturalistic clips of individuals engaging in various human-computer interaction (HCI) tasks guided by slides, recorded via a laptop camera and microphone. Its subset, the AVEC 2014 depression corpus (Valstar et al., 2014), only includes two HCI tasks: (1) FreeForm, where participants respond to predefined questions in German; and (2) NorthWind, where they read an excerpt from Die Sonne und der Wind. Unlike AVEC 2013, each task is treated as an independent clip, yielding 300 short clips. Both datasets provide clip-level depression severity annotations based on BDI-II scores. However, a key limitation is the overlap of individuals between training/validation and test sets, making the evaluation less objective. While the AVEC 2013/2014 datasets are limited in size for training and evaluation, the Vlog Depression Dataset (Yoon et al., 2022) offers a large-scale alternative. It contains 961 vlogs from 816 individuals with a balanced distribution of depressed and non-depressed examples. Unlike the above controlled datasets, these vlogs were obtained from YouTube, where individuals exhibit natural, unrestricted behaviours. However, their binary depression labels were assigned solely by non-expert annotators (college students) based on vlog content and titles, which introduces potential label reliability concerns.
@@
Alternatively, a large part of public ADA datasets were collected under clinical interview-style scenarios. The DAIC-WOZ dataset (Gratch et al., 2014) is an audio-visual-text corpus recorded under several semi-structured interviews designed to assess PTSD and depression. Participants undergo one of four interview types: face-to-face, video conference, Wizard-of-Oz (human-controlled virtual interviewer), or pre-defined virtual agent-controlled interviews. To avoid ethical risks, the dataset provides de-identified audio, text transcripts, and mid-/low-level anonymous facial behaviour features extracted by OpenFace (Baltrušaitis et al., 2016), rather than raw facial videos. Clip-level depression severity is annotated using self-reported PHQ-8 scores (PHQ-9 without the suicide question). Its extension, Extended DAIC-WOZ (Ringeval et al., 2019), used for the AVEC 2019 challenge, follows the same protocol but includes more clips (i.e., 275 clips from 275 individuals), which incorporates additional deep spectrum audio features and deep facial features extracted via VGG-16 and ResNet-50. Similarly, the Chinese Multimodal Depression Corpus (CMDC) (Zou et al., 2022) also comprises face-to-face human interviews conducted under strict participant selection criteria, with assessments by professional psychiatrists. Interviews follow pre-defined topics based on clinical practice and depression questionnaires, yielding de-identified transcripts, audio signals, and OpenFace facial primitives, where clip-level depression labels are defined by HAMD-17 and PHQ-9 scores. The Pittsburgh Clinical Depression (PCD) dataset (Dibeklioğlu et al., 2018) also records in-person face-to-face clinical interviews with participants undergoing treatment via either selective serotonin re-uptake inhibitor (SSRI) or Interpersonal Psychotherapy (IPT), where interview questions and depression categories are both defined by the HAM-D questionnaire.
@@
In addition to the above datasets recording facial, audio and/or textual behaviours, other datasets explore alternative modalities for ADA. The Looking at the Body (LaB) dataset (Lin et al., 2021) captures full-body motions, gestures, facial behaviours and speeches from 35 gender-balanced participants in face-to-face interviews with open-ended questions. Each of its clip is annotated with self-reported PHQ-8 scores, along with anxiety, stress, somatic symptom burden, and personality labels. The Emotional Audio-Textual Depression Corpus (EATD-Corpus) (Shen et al., 2022) contains audio and transcript recordings from 162 Chinese students participating in APP-based online interviews. Each participant answered three randomly selected questions, with depression severity labeled using the self-reported SDS questionnaire (Zung, 1965). The MODMA dataset (Cai et al., 2022) record EEG and speech data from 24 clinically depressed patients and 29 healthy individuals. It includes: (1) 128-electrode EEG recordings during resting state and a Dot Probe task; (2) 3-electrode EEG signals in resting state; and (3) audio recordings from tasks of 18-question interviews, word-reading, and facial expression picture description. The binary depression labels were manually assigned by professional psychiatrists. Although the lack of large-scale and well-annotated dataset challenge remains, the above datasets form a first step towards addressing this gap. Overall, these are the main datasets made publicly available and have been more commonly researched within the community.
6. Conclusion and Discussion

### v2

@@ 4.7.3 Audio-visual-language approaches
Hand-crafted approaches: Based on audio, visual and texts, Williamson et al. [268] analyse both semantic contents and context from dialogue transcripts, where contents are modeled at the sentence-level via the GloVe embedding [294] followed by PCA/ZCA to reduce dimension and sparse coding to extract final semantic features. To capture the joint dynamical facial behaviour properties during speech, a rank-ordered eigenspectrum is computed from multiple AUs, describing within-and cross-channel distributional properties of the multi-channel AU time series. Finally, the extracted audio (discussed in Sec. 4.5), visual and language features are individually fed to a Gaussian Staircase Model to make unimodal depression prediction, followed by a decision-level fusion. Building on a similar pipeline as AVEC 2013/2014 baselines, Pampouchidou et al. [25] achieve the clip-level multi-modal depression assessment via two strategies: (i) decision-level fusion of depression predictions made by clip-level visual representation (i.e., statistical and spectral features of facial landmarks, head motions and blinking rates), clip-level audio representation computed from various audio LLDs, and the clip-level language representation; and (ii) fusing multi-modal clip-level feature and then feeding it to a SVR or RF for ADA. Gupta et al. [251] extract frame-level LBP features and facial landmarks, alongside LBP-TOP and optical-flow-based motion vectors to capture short-term facial dynamics, all of which summarised as clip-level statistics. Meanwhile, multiple audio LLDs, spectral shape, spectro-temporal modulations, pitch periodicity, and long-term spectral variability, are jointly form a clip-level audio representation. Linguistic features are derived by mapping word sequences to affective ratings, processed via a Distributional Semantics Model and PCA as clip-level representation. The three modalities are then integrated and fed into an SVR for clip-level ADA. Nasir et al. [239] compute arithmetic means of: (1) frame-level pitch, glottal and voice quality-related as well as vocal tract audio (e.g., F1, F2 and F3) features; and (2) several facial behaviour primitives expressed in each 10s window. Additionally facial feature dynamics (i.e., velocity and acceleration), facial geometric and facial topology changes defined by facial landmarks are also computed. Finally, the feature-level fusion is achieved on clip-level audio, visual and language (statistics of phoneme rate and duration) representations, followed by RF for ADA.
@@
Alternatively, some multi-modal approaches learn visual depression cues from pre-extracted facial/visual behaviour primitives rather than original videos. Based on virtual assistant-guided sessions, Wu et al. [221] separately conduct naive decision-level fusion and feature-level fusion for various mid-level and low-level hand-crafted features extracted from text, audio, facial attributes, heart rate, and eye movement, where MLP is employed as the classifier for four-level depression classification. Yang et al. [275, 267] first select several equal-length video segment within every video to balance the depressed/non-depressed training examples. Based on each segment, it learns a Histogram of Displacement Range representing the range and speed of facial landmark displacements. Meanwhile, segment-level statistics summarising hand-crafted LLD audio features are computed. Finally, segment-level audio-visual features are processed by a two-stream CNN to individually make depression predictions, which are fused with textual features via a decision tree. Gong et al. [36] specifically explore the relationship between human audio, facial and language behaviours triggered by $8383$ different interview topics defined by DAIC-WOZ database [64] and the corresponding depression severity, where language information are partially used for topic modelling. For each topic, clip-level statistics of frame-level audio LLDs and facial descriptors as well as LIWC-based textual features, are extracted. Finally, a vector summarising all topic-level multi-modal features is constructed to make the clip-level depression prediction.
@@
Deep learning approaches: CNN-RNN framework have been widely extended for audio-visual-language ADA. Zhang et al. [270] deep learn frame-level/short-term audio, facial and language features via three CNNs, and then model their long-term dynamics via a BiLSTM. Then, the features learned from three modalities are fused via an attention, followed by a MLP to make the depression prediction. Wei et al. [245] first apply CNN-BiLSTM to independently learn audio, visual and textual features, and then introduce a novel sub-attention to fuse these multi-modal clip-level features. Lin et al. [262] also propose a Hierarchy BiLSTM to process audio, visual and text data for ADA. Rodrigues et al. [37] apply a GNN-LSTM to process deep spectrum audio features and deep face features provided by ResNet-50. It additionally employs BERT to extract language features from texts, followed by a CNN-LSTM model to learn clip-level language depression features. Finally, multi-modal features are concatenated and fed to a dense layer for ADA. The similar strategy was also employed by [295], where three BiLSTM networks are developed to obtain clip-level audio, visual and language representations from frame/segment-level hand-crafted and deep-learned features, followed by a multi-layer attention network to fuse them for ADA. In [273], a ResNet and Swin Transformer-based dual visual encoder first learns global and local spatial facial features from each frame, and weights each facial local region’s importance. A BiLSTM and a multi-head attention operation then model spatio-temporal depression cues from all frame-level representations. Meanwhile, the Mel spectrogram is calculated from the audio signal, followed by GRU and multi-head attention for temporal modelling. The text feature is modelled via a BERT Chinese model and attention. Finally, another attention module selects and combines depression features of three modalities, and a MLP is employed for the final prediction. Alternatively, some approaches achieve deep audio-visual-language ADA without using RNN-style networks, Haque et al. [276] extend the Causal CNN to learn sentence-level embeddings from audio, text and 3D facial landmarks, which are summarised into a single vector via simple average at the sentence-level, and then fed to SVM for ADA. Sun et al. [260] propose two audio, visual and language fusion strategies for ADA, where one forms an cubeMLP based on three independent MLP units to mix multi-modal features across three axes, while the other [261] designing a tensor-based multi-modal Transformer conducting source-target attentions between features of every modality pair. Zhang et al. [263] introduces two attention-based aggregation strategies to: (1) dynamically evaluate the relevance among audio, visual and language features to adaptively integrate their temporal dynamics; and (2) determine the importance of frame-level amalgamated features to select the most depression-specific features. To investigate dynamics of multi-modal behaviours, Shen et al. [266] propose a multi-modal behaviour graph representation learning framework which learns a set of heterogeneous node features to represent short segment-level facial behaviours, long segment-level audio behaviours, and the global text cues. Then, edges are constructed to connect nodes of the same and different modalities, to model multi-scale intra and inter-modal dynamics. This clip-level graph representation is then fed to a GNN for ADA.
4.7.4 Other multi-modal approaches
@@
A few studies also utilise other modality combination for ADA. Liu et al [198] propose a two-stream CNN for facial and body movement-based ADA, where one stream is pre-trained to predict frame-level facial expression, based on which a video-level expression entropy is computed. The other stream locates the position of the individual’s 18 body joints and calculate their video-level motion amplitude. Finally, a LR fuses video-level facial expression entropy and body joint motion amplitude for ADA. Ning et al. [242] propose an auxiliary decision-making system to combine EEG and speech signals for ADA. It explores both linear and nonlinear features from each modality, and concatenates them for depression classification. Li et al. [296] fuse a sparse emotion dictionary describing textual data and spectral audio features for ADA, where a multi-head self-attention-based low and high frequency feature fusion strategy is proposed. Ye et al. [243] extract several hand-crafted audio LLDs and apply DeepSpectrum to deep learn depression-related audio features, which are subsequently integrated into a customised TCN, with the final layer employing relational attention classification. Meanwhile, they employ the Continuous Bag of Words (CBOW) method for text feature extraction. Both clip-level features are concatenated and fed to a MLP for ADA. Zheng et al. [246] propose a temporal convolution transformer to jointly assess depression and emotion status from human audio, visual, text and additional physiological behaviours. Fan et al. [264] leverage CNN, transformer and GNN to extract depression-related frame/segment-level video, audio and remote photoplethysmographic (rPPG) features and treats each of them as a graph node. Then, a transformer is introduced to model inter-modal, intra-modal, and tri-modal relationships among nodes learned from three modalities. Finally, a GNN fuses all features for clip-level multi-modal ADA. Lin et al. [67] extract various hand-crafted frame-level features from human body movements, facial behaviours and audio signals, which are summarised as a multi-modal clip-level representation through a Multi-modal Deep Denoising Auto-Encoder and Fisher vector encoding, followed by a RF for feature selection. Finally, MLP and LR are separately applied for ADA. Shao et al. [247] propose to learn complementary depression-related cues from brain behaviours detected by EEG and fNIRS, where spatio-temporal CNNs and transformers are stacked to jointly process both modalities.
4.8 Statistics and trend
Fig. 4 illustrates statistics computed from ADA publications considered by this survey. It can be observed that the number of ADA publications, especially for face, text and multi-modal based approaches, has drawn clear attentions since 2013. This may largely contributed by the AVEC 2013 and AVEC 2014 challenges that provided initial venues and open-source audio-visual ADA datasets for developing and comparing audio-visual ADA solutions. The small drop in 2019 might caused by the fact that there was no depression sub-challenge in AVEC 2018. Importantly, the number of ADA publications has been largely increased since 2018, suggesting that the ADA research has drawn more attentions in recent years, even during and after the COVID pandemic. Among all surveyed ADA solutions (since 2012), deep learning models have been applied to more than half of them. From another perspective, the most popular uni-modality for the surveyed ADA methods is the facial behaviour. Although a large number of audio and text-based depression feature learning strategies have been investigated, they were generally combined or even integrated with the visual modality to form multi-modal ADA methods. Authors affiliated with institutions in China, USA, and UK developed most of existing ADA solutions, with the remainder mainly coming from Institutions of other Europe and Asia countries (i.e., only counting first and last authors of each publication).
Figure 4: Statistics and trend of ADA approaches (2012-2024) surveyed in this paper.
5 Datasets and Competitions
5.1 Publicly available ADA datasets
@@
This section reviews existing publicly available ADA datasets. The main challenge is their limited scale in participant numbers and clip lengths, i.e., they contain (except the V-log dataset that lacks rigorous labeling) fewer than 200 participants with each clip usually lasting less than 30 mins (i.e., depression is a long-term condition reflected over extended periods). Another limitation is that depression labels are typically obtained from self-reported questionnaires rather than standardised clinical assessments, leading to some degrees of subjectivity and inconsistency. Such issues not only limit the advancement of DL-based ADA models but also undermine evaluation objectivity.
@@
The AVEC 2013 depression corpus [59] is a pioneering publicly available audio-visual ADA dataset containing 150 naturalistic clips of individuals engaging in various human-computer interaction (HCI) tasks guided by slides, recorded via a laptop camera and microphone. Its subset, the AVEC 2014 depression corpus [60], only includes two HCI tasks: (1) FreeForm, where participants respond to predefined questions in German; and (2) NorthWind, where they read an excerpt from Die Sonne und der Wind. Unlike AVEC 2013, each task is treated as an independent clip, yielding 300 short clips. Both datasets provide clip-level depression severity annotations based on BDI-II scores. However, a key limitation is the overlap of individuals between training/validation and test sets, making the evaluation less objective. While the AVEC 2013/2014 datasets are limited in size for training and evaluation, the Vlog Depression Dataset [68] offers a large-scale alternative. It contains 961 vlogs from 816 individuals with a balanced distribution of depressed and non-depressed examples. Unlike the above controlled datasets, these vlogs were obtained from YouTube, where individuals exhibit natural, unrestricted behaviours. However, their binary depression labels were assigned solely by non-expert annotators (college students) based on vlog content and titles, which introduces potential label reliability concerns.
@@
Alternatively, a large part of public ADA datasets were collected under clinical interview-style scenarios. The DAIC-WOZ dataset [64] is an audio-visual-text corpus recorded under several semi-structured interviews designed to assess PTSD and depression. Participants undergo one of four interview types: face-to-face, video conference, Wizard-of-Oz (human-controlled virtual interviewer), or pre-defined virtual agent-controlled interviews. To avoid ethical risks, the dataset provides de-identified audio, text transcripts, and mid-/low-level anonymous facial behaviour features extracted by OpenFace [297], rather than raw facial videos. Clip-level depression severity is annotated using self-reported PHQ-8 scores (PHQ-9 without the suicide question). Its extension, Extended DAIC-WOZ [63], used for the AVEC 2019 challenge, follows the same protocol but includes more clips (i.e., 275 clips from 275 individuals), which incorporates additional deep spectrum audio features and deep facial features extracted via VGG-16 and ResNet-50. Similarly, the Chinese Multimodal Depression Corpus (CMDC) [65] also comprises face-to-face human interviews conducted under strict participant selection criteria, with assessments by professional psychiatrists. Interviews follow pre-defined topics based on clinical practice and depression questionnaires, yielding de-identified transcripts, audio signals, and OpenFace facial primitives, where clip-level depression labels are defined by HAMD-17 and PHQ-9 scores. The Pittsburgh Clinical Depression (PCD) dataset [66] also records in-person face-to-face clinical interviews with participants undergoing treatment via either selective serotonin re-uptake inhibitor (SSRI) or Interpersonal Psychotherapy (IPT), where interview questions and depression categories are both defined by the HAM-D questionnaire.
@@
In addition to the above datasets recording facial, audio and/or textual behaviours, other datasets explore alternative modalities for ADA. The Looking at the Body (LaB) dataset [67] captures full-body motions, gestures, facial behaviours and speeches from 35 gender-balanced participants in face-to-face interviews with open-ended questions. Each of its clip is annotated with self-reported PHQ-8 scores, along with anxiety, stress, somatic symptom burden, and personality labels. The Emotional Audio-Textual Depression Corpus (EATD-Corpus) [69] contains audio and transcript recordings from 162 Chinese students participating in APP-based online interviews. Each participant answered three randomly selected questions, with depression severity labeled using the self-reported SDS questionnaire [298]. The MODMA dataset [70] record EEG and speech data from 24 clinically depressed patients and 29 healthy individuals. It includes: (1) 128-electrode EEG recordings during resting state and a Dot Probe task; (2) 3-electrode EEG signals in resting state; and (3) audio recordings from tasks of 18-question interviews, word-reading, and facial expression picture description. The binary depression labels were manually assigned by professional psychiatrists. Although the lack of large-scale and well-annotated dataset challenge remains, the above datasets form a first step towards addressing this gap. Overall, these are the main datasets made publicly available and have been more commonly researched within the community.
5.2 Grand challenges (Competitions)
To provide venues for researchers showing their ADA solutions and exchanging ideas, a series of AVEC depression challenges have been organized, which mainly focused on inferring depression via non-invasive ways, i.e., through non-verbal audio-visual spatio-temporal behaviours and verbal textual cues. Specifically, the AVEC 2013 depression challenge [59] is the first public audio-visual ADA challenge facilitating researchers from various institutions to extend the analysis of affective behaviours to infer a more complex mental state: depression severity, which brought together the initial efforts for audio-visual ADA. Its subsequent AVEC 2014 challenge [60] follows the same manner but only used two task-dependent subsets (i.e., AVEC 2014 depression dataset). In both challenges, participant teams were required to develop their models based on original face videos and German speech signals. The ADA baselines of both challenges extract various hand-crafted audio and visual features while leveraging several traditional ML regressors for clip-level depression severity estimation (please check the main paper for details). The Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) were employed as metrics for the evaluation.
The following AVEC 2016 [61], AVEC 2017 [62] and the most recent AVEC 2019 depression [63] challenges focused on de-identified multi-modal human behaviour-based ADA, which additionally introduced verbal text modality. They employed DAIZ-WOZ depression dataset [64] or its extension to provide a common ADA benchmark under well-defined and strictly comparable semi-structured clinical interview-based conditions. In these challenges, only de-identified audio and text data as well as anonymous facial behaviour primitives/attributes (e.g., AUs, facial landmarks, head poses, gazes, HOG features and discrete emotions) were provided. Participant teams were required to develop their ADA system based on such de-identified behaviour data rather than original audio signals and face videos. Besides the depression severity estimation, these challenges also provide baselines for binary depression classification. Again, the RMSE and MAE were used for evaluating depression severity estimation systems, while F1, Precision and Recall were adopted to evaluate the depression classification systems. Additionally, the Concordance Correlation Coefficient (CCC) that takes both Mean Square Error (MSE) and Pearson Correlation Coefficient (PCC) between depression severity predictions and their ground-truths into consideration was used to rank all participant teams in AVEC 2019.
6 Conclusion and Discussion